What volume (in mL) of 0.844 M NaOH solution is required to react exactly with 48.5 mL of 1.39 M H2SO4 solution?
Answer #1

Balanced chemical equation is:

H2SO4 + 2 NaOH ---> Na2SO4 + 2 H2O

Here:

M(H2SO4)=1.39 M

M(NaOH)=1.39 M

V(H2SO4)=48.5 mL

According to balanced reaction:

2*number of mol of H2SO4 =1*number of mol of NaOH

2*M(H2SO4)*V(H2SO4) =1*M(NaOH)*V(NaOH)

2*1.39 M *48.5 mL = 1*1.39M *V(NaOH)

V(NaOH) =

Answer: 97.0 mL